# Superamalgamation for modal lattices via non-distributive dualities

### v1

@@ 3.2 Craig interpolation for weak positive logic
The fact that ${\uparrow}f_{1}[U]$ is a filter is easy to see. To show that the complement of ${\downarrow}f_{2}[Y_{2}\setminus V]$ is a filter, take $x,y\notin{\downarrow}f_{2}[Y_{2}\setminus V]$. Suppose that $x\curlywedge y\preceq f_{2}(z)$ where $z\notin V$. Because $f_{2}$ is an L-morphism, there are $x^{\prime},y^{\prime}$ such that $x^{\prime}\curlywedge y^{\prime}\preceq z$, $x\preceq f_{2}(x^{\prime})$ and $y\preceq f_{2}(y^{\prime})$. Because $x\notin{\downarrow}f_{2}[Y_{2}\setminus V]$, we have $x^{\prime}\in V$, for otherwise, $x^{\prime}\in Y_{2}\setminus V$ so $x\preceq f_{2}(x^{\prime})$ gets us a contradiction. Similarly, $y^{\prime}\in V$, so because $V$ is a filter, $x^{\prime}\curlywedge y^{\prime}\in V$. But $V$ is a filter, so $z\in V$, a contradiction.
Both sets are closed. Indeed, $f_{2}[Y_{2}\setminus V]$ and $f_{1}[U]$ are compact, as continuous image of a compact set, and therefore are closed. By 3.9, the sets ${\downarrow}f_{2}[Y_{2}\setminus V]$ and ${\uparrow}f_{1}[U]$ are closed.
For the second statement, assume that $y\in{\uparrow}f_{1}[U]\cap{\downarrow}f_{2}[Y_{2}\setminus V]$. Then there is a point $x\in U$ and $z\in Y_{2}\setminus V$ such that $f_{1}(x)\preceq y\preceq f_{2}(z)$. Now because $f_{2}$ is surjective, there is some $t$ such that $f_{1}(x)=f_{2}(t)$. Then $f_{2}(z)\curlywedge f_{2}(t)=f_{1}(x)$, so $f_{2}(z\curlywedge t)=f_{1}(x)$. Hence $(x,z\curlywedge t)\in\mathrm{Pb}(f_{1},f_{2})$, and this belongs to $\pi_{1}^{-1}[U]$. By assumption, this also belongs to $\pi_{2}^{-1}[V]$, so $z\curlywedge t\in V$, and by upwards closure, $z\in V$, a contradiction. This shows the result. ∎
Now assume then that $U\subseteq Y_{1}$ and $V\subseteq Y_{2}$ are clopen filters, such that $\pi_{1}^{-1}[U]\subseteq\pi_{2}^{-1}[V]$. By Claim 3.13, we get that there is a clopen filter $W$ such that ${\uparrow}f_{1}[U]\subseteq W$ and ${\downarrow}f_{2}[Y_{2}\setminus V]\subseteq X\setminus W$. This then implies that $U\subseteq f_{1}^{-1}[W]$ and $f_{2}^{-1}[W]\subseteq V$, which proves superamalgamation. ∎
@@ 4.1 Duality for modal lattices
In this section we look at superamalgamation of modal lattices, using the general strategy employed in the previous section. For that purpose, we extend the duality used in the previous section. Let us first introduce the modal version of L-frames, first introduced in [1].
@@ 4.2 Craig interpolation for weak positive modal logic
The last two conditions are easy, as everything is defined componentwise. Now let us show that the projections $\pi_{i}$ are bounded L-morphisms. We do it for $\pi_{1}$, showing that it satisfies all three conditions.
@@
    This one is obvious, by definition of $R$ on $\mathrm{Pb}(f_{1},f_{2})$.
@@
    Assume that $\pi_{1}(x_{1},x_{2})\mathrel{R}z_{1}$, that is, that $x_{1}\mathrel{R}z_{1}$. Let $x\coloneq f_{1}(x_{1})=f_{2}(x_{2})$ and $z\coloneq f_{1}(z_{1})$. Then $x\mathrel{R}z$, and we can find a $t\preceq z$ that is minimal in $R[x]$. As $f_{1}(x_{1})\mathrel{R}t$, we can find $y_{1}\in Y_{1}$ and $y_{2}\in Y_{2}$ such that $x_{1}\mathrel{R}y_{1}$, $x_{1}\mathrel{R}y_{2}$, $f_{1}(y_{1})\preceq t$ and $f_{2}(y_{2})\preceq t$. But then, we have $f_{1}(y_{1}),f_{2}(y_{2})\in R[x]$, and as $t$ is minimal, $f_{1}(y_{1})=t=f_{2}(y_{2})$. Thus the pair $(y_{1},y_{2})$ is in $\mathrm{Pb}(f_{1},f_{2})$, and we have $(x_{1},x_{2})\mathrel{R}(y_{1},y_{2})$ and $\pi_{1}(y_{1},y_{2})\preceq z$.
@@
    We proceed similarly to the previous condition. ∎
Combining this with Theorem 2.17, we get the following.
@@ 4.3 Craig interpolation in extensions
Let us start by introducing the L-frame semantics for our logic.
@@
One should observe that in the classical case, the directed modal logic $\mathsf{K.2}$ doesn’t have Craig interpolation. This makes the above result quite surprising. This can be explained by the fact that when working with modal L-spaces, many quantifiers can be eliminated by picking suitable maximal or minimal elements: in short, the usual impediments to amalgamation deriving from the need for choices disappear, since these choices become canonical.

### v2

@@ 3.2 Craig interpolation for weak positive logic
The fact that ${\uparrow}f_{1}[U]$ is a filter is easy to see. To show that the complement of ${\downarrow}f_{2}[Y_{2}\setminus V]$ is a filter, take $x,y\notin{\downarrow}f_{2}[Y_{2}\setminus V]$. Suppose that $x\curlywedge y\preceq f_{2}(z)$ where $z\notin V$. Because $f_{2}$ is an L-morphism, there are $x^{\prime},y^{\prime}$ such that $x^{\prime}\curlywedge y^{\prime}\preceq z$, $x\preceq f_{2}(x^{\prime})$ and $y\preceq f_{2}(y^{\prime})$. Because $x\notin{\downarrow}f_{2}[Y_{2}\setminus V]$, we have $x^{\prime}\in V$, for otherwise, $x^{\prime}\in Y_{2}\setminus V$ so $x\preceq f_{2}(x^{\prime})$ gets us a contradiction. Similarly, $y^{\prime}\in V$ and since $V$ is a filter, $x^{\prime}\curlywedge y^{\prime}\in V$. But $V$ is a filter, so $z\in V$, a contradiction.
Both sets are closed. Indeed, $f_{2}[Y_{2}\setminus V]$ and $f_{1}[U]$ are compact, as they are images of a compact set via a continuous map, and therefore are closed. By Proposition 3.9, the sets ${\downarrow}f_{2}[Y_{2}\setminus V]$ and ${\uparrow}f_{1}[U]$ are closed.
For the second statement, assume that $y\in{\uparrow}f_{1}[U]\cap{\downarrow}f_{2}[Y_{2}\setminus V]$. Then there are points $x\in U$ and $z\in Y_{2}\setminus V$ such that $f_{1}(x)\preceq y\preceq f_{2}(z)$. Now since $f_{2}$ is surjective, there is some $t$ such that $f_{1}(x)=f_{2}(t)$. Then $f_{2}(z)\curlywedge f_{2}(t)=f_{1}(x)$ and thus $f_{2}(z\curlywedge t)=f_{1}(x)$. Hence $(x,z\curlywedge t)\in\mathrm{Pb}(f_{1},f_{2})$, and this belongs to $\pi_{1}^{-1}[U]$. By assumption, this also belongs to $\pi_{2}^{-1}[V]$, so $z\curlywedge t\in V$, and by upwards closure, $z\in V$, a contradiction. This proves the result. ∎
Now assume that $U\subseteq Y_{1}$ and $V\subseteq Y_{2}$ are clopen filters, such that $\pi_{1}^{-1}[U]\subseteq\pi_{2}^{-1}[V]$. By Claim 3.13, we obtain that there is a clopen filter $W$ such that ${\uparrow}f_{1}[U]\subseteq W$ and ${\downarrow}f_{2}[Y_{2}\setminus V]\subseteq X\setminus W$. This implies that $U\subseteq f_{1}^{-1}[W]$ and $f_{2}^{-1}[W]\subseteq V$, which proves superamalgamation. ∎
@@ 4.1 Duality for modal lattices
In this section we look at superamalgamation of modal lattices, using the general strategy employed in the previous section. For that purpose, we extend the duality used in the previous section. Let us first recall the modal version of L-frames, first introduced in [1].
@@ 4.2 Craig interpolation for weak positive modal logic
The last two conditions are easy to verify, since everything is defined componentwise. Now we show that the projections $\pi_{i}$ are bounded L-morphisms. We do it for $\pi_{1}$, showing that it satisfies all three conditions.
@@
    Follows straightforwardly from the definition of $R$ on $\mathrm{Pb}(f_{1},f_{2})$.
@@
    Assume that $\pi_{1}(x_{1},x_{2})\mathrel{R}z_{1}$, that is, $x_{1}\mathrel{R}z_{1}$. Let $x\coloneq f_{1}(x_{1})=f_{2}(x_{2})$ and $z\coloneq f_{1}(z_{1})$. Then $x\mathrel{R}z$, and we can find a $t\preceq z$ that is minimal in $R[x]$. As $f_{1}(x_{1})\mathrel{R}t$, we can find $y_{1}\in Y_{1}$ and $y_{2}\in Y_{2}$ such that $x_{1}\mathrel{R}y_{1}$, $x_{1}\mathrel{R}y_{2}$, $f_{1}(y_{1})\preceq t$ and $f_{2}(y_{2})\preceq t$. But then, we have $f_{1}(y_{1}),f_{2}(y_{2})\in R[x]$, and as $t$ is minimal, $f_{1}(y_{1})=t=f_{2}(y_{2})$. Thus the pair $(y_{1},y_{2})$ is in $\mathrm{Pb}(f_{1},f_{2})$, and we have $(x_{1},x_{2})\mathrel{R}(y_{1},y_{2})$ and $\pi_{1}(y_{1},y_{2})\preceq z$.
@@
    We proceed similarly to the previous case. ∎
Combining this with Theorem 2.17, we obtain the following.
@@ 4.3 Craig interpolation in extensions
We start by recalling from [1] the L-frame semantics for our logic.
@@
One should observe that in the classical case, the directed modal logic $\mathsf{K.2}$ does not have Craig interpolation [20, Section 5.6.2]. This makes the above result quite surprising. This can be explained by the fact that when working with modal L-spaces, many quantifiers can be eliminated by picking suitable maximal or minimal elements: in short, the usual impediments to amalgamation deriving from the need for choices disappear, since these choices become canonical.